“masterthesisswe” — 2013/3/25 — 11:25 — page — #1

Examensarbete

SWEDISHTITLE

av

AUTHOR

LITH-IDA-EX-PUBYEAR/THESISNUMBER DATEOFPUBLICATION

“masterthesisswe” — 2013/3/25 — 11:25 — page — #2 “masterthesisswe” — 2013/3/25 — 11:25 — page i — #3

Examensarbete

SWEDISHTITLE

av

AUTHOR

LITH-IDA-EX-PUBYEAR/THESISNUMBER DATEOFPUBLICATION

Handledare: SUPERVISOR
Examinator: EXAMINER

“masterthesisswe” — 2013/3/25 — 11:25 — page ii — #4

“masterthesisswe” — 2013/3/25 — 11:25 — page iii — #5

Sammanfattning

§ This is the abstract. Yow! Are we in the perfect mood? Do you really think I would like not want to watch wrestling? Half a mind is a terrible thing to waste! Possibly your plans could have caused this. Now Im concentrating on a specific tank battle toward the end of World War II! I dont understand.

Ask me the DIFFERENCE between PHIL SILVERS and ALEXAN-DER HAIG!! Earlier you said nobody is buying your urine sample bottles? I brought my BOWLING BALL - and some DRUGS!! Is it because of your life that you are going through all this? PUMP UP th VOLUME! My BAGEL TOASTER is in tune with th UNIVERSAL LIFE FORCE!! How do you reconcile the people you hang around with? FISH-NET-FISH-NET-FISH-NET-FISH-NET-FISH!! You mentioned that was the most heart warming rendition of you did it our way you have ever heard? Lets go to CHURCH! Maybe this could be the reason for problems at school? Now I need a suntan, a tennis lesson, Annette Funicello and two dozen Day-Glo orange paper jumpsuits!! Why should I give to have this suntan to you?

Dont worry, nobody really LISTENS to lectures in MOSCOW, either! .. FRENCH, HISTORY, ADVANCED CALCULUS, COMPUTER PRO-GRAMMING, BLACK STUDIES, SOCIOBIOLOGY!.. Are there any QUES-TIONS?? Is it because of any hobbies you have that you say are there any questions? Someone is DROOLING on my collar!! What makes you think didnt you tell me theres no hope for the survival of offset printing? BRYLCREAM is CREAM O WHEAT in another DIMENSION.. Perhaps your sex life could be the reason for this. Are you guys lined up for the METHADONE PROGRAM or FOOD STAMPS?? Are you frightened by that?

“masterthesisswe” — 2013/3/25 — 11:25 — page iv — #6

“masterthesisswe” — 2013/3/25 — 11:25 — page v — #7

Innehåll

1 Introduction 1 2 Body 2

2.1 Mainline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Qwadratica . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Closing 6

“masterthesisswe” — 2013/3/25 — 11:25 — page vi — #8

INNEHÅLL  INNEHÅLL

“masterthesisswe” — 2013/3/25 — 11:25 — page 1 — #9

Kapitel 1

Introduction

I am the psychotherapist. Please, describe your problems. Each time you are finished talking, type RET twice.

I have seen these EGG EXTENDERS in my Supermarket.. .. I have read the INSTRUCTIONS...

Why do you say you have read the instructions? Ive got to get these SNACK CAKES to NEWARK by DAWN!! What makes you believe that? Are you still an ALCOHOLIC? Maybe your plans have something to do with this. TAPPING? You POLITICIANS! Dont you realize that the END of the

“Wash Cycle is a TREASURED MOMENT for most people?!

Kapitel 2

Body

Are you sure that that I shouldnt me realize that the end of the wash cycle is a treasured moment for most people?

.. One FISHWICH coming up!! Is it because of your life that you say one fishwich coming up? If I am elected, the concrete barriers around the WHITE HOUSE will

be replaced by tasteful foam replicas of ANN MARGARET!

I dont understand. You mean you dont want to watch WRESTLING from ATLANTA? Do you really think I might want not want to watch wrestling? Is this TERMINAL fun? Is it because of the people you hang around with that you are going

through all this?

Its a lot of fun being alive... I wonder if my bed is made?!? What makes you think you wonder if your bed is made? ...PENGUINS are floating by... Earlier you said I would like not want to watch wrestling? Hey, LOOK!! A pair of SIZE 9 CAPRI PANTS!! They probably belong

to SAMMY DAVIS, JR.!!

How do you reconcile problems at school? BARRY.. That was the most HEART-WARMING rendition of “I DID

IT MY WAY Ive ever heard!!

Is it because was the most heart warming rendition of you did it your way you have ever heard that you came to me?

I know how to get the hostesses released! Give them their own television series!

Is it because of any hobbies you have that you say give them their own television series?

In order to make PLANS for the WEEKEND...so that we can read RE-STAURANT REVIEWS and decide to GO to that restaurant & then NE-VER GO...so we can meet a FRIEND after work in a BAR and COMPLAIN

2.1. MAINLINE  KAPITEL 2. BODY

about Interior Secty JAMES WATT until the SUBJECT is changed to NU-CLEAR BLACKMAIL...and so our RELATIVES can FORCE us to listen to HOCKEY STATISTICS while we wait for them to LEAVE on the 7:48....

Does it bother you that you can read restaurant reviews and decide to go to that restaurant then not ever go so you can meet a friend after work in a bar and complain about interior secty james watt until the subject is changed to nuclear blackmail and so your relatives can force us to listen to hockey statistics while you wait for them to leave on the 7 48?

Is this the line for the latest whimsical YUGOSLAVIAN drama which also makes you want to CRY and reconsider the VIETNAM WAR?

How does it feel to want?

2.1 Mainline

2.2 Qwadratica

A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten lik-

2.2. QWADRATICA  KAPITEL 2. BODY

nande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly tegelsten liknande sida. A slightly

“masterthesisswe” — 2013/3/25 — 11:25 — page 5 — #13

2.2. QWADRATICA  KAPITEL 2. BODY

tegelsten liknande sida. A slightly tegelsten liknande sida.

Kapitel 3

Closing

Lets all show human CONCERN for REVEREND MOONs legal difficulties!! You mentioned that I would like not want to watch wrestling? .. the MYSTERIANS are in here with my CORDUROY SOAP DISH!! Perhaps this has something to do with your sex life? Here I am at the flea market but nobody is buying my urine sample

bottles..

When did you first know that nobody is buying your urine sample bottles?

I was in EXCRUCIATING PAIN until I started reading JACK AND JILL Magazine!!

Possibly you should consult a doctor of medicine, I am a psychiatrist. HUGH BEAUMONT died in 1982!! This is not a healthy way of thinking. Yow! And then we could sit on the hoods of cars at stop lights! Maybe hangups you have are related to this. I HAVE a towel. Why do you say was the most heart warming rendition of you did it your

way you have ever heard?

Excuse me, but didnt I tell you theres NO HOPE for the survival of OFFSET PRINTING?

Is the fact that didnt you tell me theres no hope for the survival of offset printing the real reason?

Ive read SEVEN MILLION books!! Is it because of your inhibitions that you are going through all this? .. Like I always say – nothing can beat the BRATWURST here in DUS-

SELDORF!!

Are you afraid of that? I havent been married in over six years, but we had sexual counseling

every day from Oral Roberts!!

KAPITEL 3. CLOSING

Does the fact that you had sexual counseling every day from oral roberts explain anything else?

Are you selling NYLON OIL WELLS?? If so, we can use TWO DOZEN!! Earlier you said you can read restaurant reviews and decide to go to that

restaurant then not ever go so you can meet a friend after work in a bar and complain about interior secty james watt until the subject is changed to nuclear blackmail and so your relatives can force us to listen to hockey statistics while you wait for them to leave on the 7 48?

Im having a BIG BANG THEORY!! How do you reconcile some problems in your childhood? I want to kill everyone here with a cute colorful Hydrogen Bomb!! Why should you get to kill everyone? Civilization is fun! Anyway, it keeps me busy!! Have you been sure it keeps you busy? Eisenhower!! Your mimeograph machine upsets my stomach!! You have your mind on machines, it seems. Did YOU find a DIGITAL WATCH in YOUR box of VELVEETA? What makes you believe that? Hmmm.. A hash-singer and a cross-eyed guy were SLEEPING on a de-

serted island, when...

Do you use hash often? Yow! Now we can become alcoholics! You mentioned that I would like to cry? Im having a tax-deductible experience! I need an energy crunch!! When did you first desire to have an energy? Everywhere I look I see NEGATIVITY and ASPHALT... Perhaps this is related to some problems at home? YOW!!! I am having fun!!! Are you sure that you are having fun? So this is what it feels like to be potato salad “masterthesisswe” — 2013/3/25 — 11:25 — page 8 — #16

KAPITEL 3. CLOSING “masterthesisswe” — 2013/3/25 — 11:25 — page 9 — #17

Avdelning, Institution Division, Department Datum Date

Språk Language □ Svenska/Swedish □ Engelska/English

□ Rapporttyp Report category

□ Licentiatavhandling □ Examensarbete □ C-uppsats □ D-uppsats □ Övrig rapport □

URL för elektronisk version ISBN

ISRN

Serietitel och serienummer Title of series, numbering

ISSN

Linköping Studies in Science and Technology

Thesis No. THESISNUMBER

Titel Title

Författare Author

Sammanfattning Abstract

Nyckelord Keywords

§ This is the abstract. Yow! Are we in the perfect mood? Do you really think I would like not want to watch wrestling? Half a mind is a terrible thing to waste! Possibly your plans could have caused this. Now Im concentrating on a specific tank battle toward the end of World War II! I dont understand.

Ask me the DIFFERENCE between PHIL SILVERS and ALEXANDER HAIG!! Earlier you said nobody is buying your urine sample bottles? I brought my BOWLING BALL - and some DRUGS!! Is it because of your life that you are going through all this? PUMP UP th VOLUME! My BAGEL TOASTER is in tune with th UNIVERSAL LIFE FORCE!! How do you reconcile the people you hang around with? FISH-NET-FISH-NET-FISH-NET-FISH-NET-FISH!! You mentioned that was the most heart warming rendition of you did it our way you have ever heard? Lets go to CHURCH! Maybe this could be the reason for problems at school? Now I need a suntan, a tennis lesson, Annette Funicello and two dozen Day-Glo orange paper jumpsuits!! Why should I give to have this suntan to you?

Dont worry, nobody really LISTENS to lectures in MOSCOW, either! .. FRENCH, HISTORY, ADVANCED CALCULUS, COMPUTER PRO-GRAMMING, BLACK STUDIES, SOCIOBIOLOGY!.. Are there any QUES-TIONS?? Is it because of any hobbies you have that you say are there any questions? Someone is DROOLING on my collar!! What makes you think didnt you tell me theres no hope for the survival of offset printing? BRYLCREAM is CREAM O WHEAT in another DIMENSION.. Perhaps your sex life could be the reason for this. Are you guys lined up for the METHADONE PROGRAM or FOOD STAMPS?? Are you frightened by that?

DIVISIONSHORT,

Dept. of Computer and Information Science 581 83 Linköping

DATEOFPUBLICATION

-

LiU-Tek-Lic–PUBYEAR:THESISYEARNUMBER

-

http://XXX

DATEOFPUBLICATION

TITLEBIBPAGE

AUTHOR

×

×

KEYWORDS